### v1

@@ 5.4 LIMITATIONS
As Bluesky is still in development, user engagement may not be as consistent as on established platforms like Twitter/X. Consequently, usage patterns might be unstable and subject to changes due to platform updates, such as the end of a public open beta, or external factors. Additionally, users on Bluesky tend to focus on topics like art and AI, which may not generalize our findings to other social media platforms. Among many social interactions on Bluesky, we focused on the networking behavior among users, excluding activities related to posts, such as replying, liking, and reposting.
5.5 FUTURE WORK
We will examine the evolutionary aspects of rumor and misinformation spread [13, 12], the role of blocking mechanisms [9, 11], and platform-specific characteristics that promote sustainability of social media platforms [15, 14]. Testing the causal effects of user attributes and behaviors, along with evaluating the predictive power of social networks, is crucial for modeling future user behavior [33, 22]. Lastly, comparing social networks across diverse online environments like Epinions, Slashdot, and Wikipedia could reveal persistent interaction patterns in user behaviors that extend beyond typical social media platforms [1, 7, 22]
6 Source Code and Scripts
@@
The scripts and dataset used for crawling Bluesky will be publicly available in our DOI [^6] https://ieee-dataport.org/documents/bluetempnet-temporal-multi-network-dataset-social-interactions-bluesky-social. The proposed graphs are saved in GEXF [^7] https://gephi.org/users/supported-graph-formats/ format and named as follows:
@@
  • graph_dimension1.gexf: This file is the Feed member interaction network, where an edge has attributes sign and time and a node represents Feed member.
@@
  • graph_dimension2.gexf: This file is the Feed creator interaction network, where an edge has attributes sign and time, and a node represents Feed creator.
@@
  • graph_dimension3.gexf: This file contains the community interaction network. Each node has a type attribute, which can be one of the following: member, creator, or Feed. Each edge includes a time attribute.
  • multi_graph.gexf: This file is the multi-network integrating the three-dimensional graphs. For the sake of simplicity, undirected edges from the community interaction network are converted to bidirectional edges.
Each timestamp in the edge data is formatted in ISO 8601 (e.g., 2023-09-22T09:32:17.974Z) based on UTC.
Data collection scripts are written in Python 3 using NetworkX version 2.6.3. These scripts are named according to each dataset (e.g., graph_dimension1.py). While user IDs remain anonymous, we provide some public user activities and information on the collected Feeds as follows:
@@
      URI (a string of characters used to uniquely identify a Feed on the internet)
@@
      Mapping between URI and node ID
@@
      Creator of Feed
@@
This work was funded by Office of Naval Research, under Award No. N00014-21-1-4002. Opinions, interpretations, conclusions, and recommendations within this article are solely those of the authors.

### v2

@@ DISCUSSION
Bluesky presents a unique opportunity for researchers to examine user behavior on social media, but it also faces significant challenges to its long-term viability. As the platform is still developing, user engagement may be less consistent than on established platforms like Twitter/X, resulting in unstable usage patterns influenced by updates and external factors. Users tend to focus on niche topics like art, which may limit the generalizability of our findings to other social media platforms. Our analysis specifically targeted networking behaviors, excluding interactions related to posts, such as replying, liking, and reposting.
Furthermore, previous research highlights that external factors, such as interactions across different social media platforms, can affect user retention on Bluesky [15]. A key issue is that Bluesky is often viewed as a complementary platform to Twitter/X rather than a direct competitor. This perception influences user preferences, as many migrating users demonstrate strong loyalty to their original platforms, particularly Twitter/X. Such platform inertia poses a significant barrier to Bluesky’s ability to retain and engage users.
In future research, we will explore the evolutionary dynamics of sentiments and rumors on the platform [14, 13], focusing on how sentiment evolves as information spreads over time [20]. Additionally, we will investigate the effects of specific networking behaviors, such as user-to-user blocking [11, 12], on the platform’s resilience during periods of significant change [16, 15]. Our study will evaluate how blocking influences social networking dynamics and platform stability, particularly as new users join. Measuring the global structural balance of social interactions, along with the causal mechanisms behind these networking patterns, can provide valuable insights into how the platform’s environment affects user behavior [8]. Last, we aim to identify consistent user interaction patterns across a variety of online social networks, including those with and without recommendation systems [1, 23].
SOURCE CODE AND SCRIPTS
@@
The scripts and dataset used for crawling Bluesky will be publicly available in the provided DOI repository. The proposed graphs are stored in GEXF [^6] https://gephi.org/users/supported-graph-formats/ as follows:
@@
  • graph_dimension1.gexf: Feed member interaction network saved in DiGraph object, where an edge has attributes sign and time and a node is Feed member.
@@
  • graph_dimension2.gexf: Feed creator interaction network saved in DiGraph object, where an edge has attributes sign and time, and a node is Feed creator.
@@
  • graph_dimension3.gexf: Community interaction network saved in a Graph object. Each node has a node attribute that can be one of the following: member, creator, or feed. Each edge has an edge attribute, either join or create, along with a time attribute.
  • multi_graph.gexf: This is a MultiGraph object that integrates the three network dimensions. To facilitate ease of use, all undirected edges in the multigraph have been converted to bidirectional edges.
Our data collection scripts are written in Python 3. These scripts utilize the NetworkX library (version $2.6.3$) and specifically the networkx.read_gexf function [^7] https://networkx.org/documentation/stable/reference/readwrite/generated/networkx.readwrite.gexf.read_gexf.html. This function helps processing different graph objects, including DiGraph, Graph, and MultiGraph. The edge data contains timestamps formatted in the ISO 8601 standard (e.g., 2023-09-22T09:32:17.974Z), all recorded in UTC. Our resources are publicly available at our repository [^8] https://github.com/ujeong1/BlueTempNet-IEEE-DATA-2024.
While user IDs and their specific profile information remain anonymous, we provide metadata that includes non-identifiable information about users and details regarding the collected Feeds, as outlined below:
@@
      Node Index (consistent across all GEXF files)
    2. 2.
@@
      Node Index (consistent across all GEXF files)
@@
      Feed URI (a unique identifier for profiles specific to the Bluesky Feed)
@@
      Creator of Feed (given as Anonymized ID)
@@
Table 8: Metadata formats in user_metadata.csv. The abbreviation b64str refers to a binary string encoded in Base64 format.
Metadata Format Example
Node Index integer 1
Anonymized ID b64str b‘gAAAAAB …ba66QOmV4’
#Followers integer 23
#Following integer 41
#Posts integer 10
Table 9: Metadata formats in feed_metadata.csv. The abbreviation b64str refers to a binary string encoded in Base64 format.
Metadata Format Example
Node Index integer 3
Feed URI string did:plc:….feed.generator/philosophy
Display Name string philosophy
Description string Feed for philosophy on Bluesky! …
Creator b64str b‘gAAAAAB . . . ba6QOmV4’
#Likes integer 10
Metadata formats are detailed in Tables 8 and 9. For clarity, examples are abbreviated with ellipses. Each row in the metadata CSV files features a unique node index alongside a user ID or Feed URI. The node index remains consistent across graph dimensions, facilitating searching nodes. A user can appear in multiple rows as both a creator and a member; for instance, the anonymized user ID b‘gAAAAAB …ba66QOmV4’ may correspond to both node index 1 and node index 2. In contrast, each node index uniquely identifies a single feed; for example, node index 3 is exclusively linked to the Feed URI did:plc:….feed.generator/philosophy.
@@
This work was funded by Office of Naval Research, under Award No. N00014-21-1-4002. Opinions, interpretations, conclusions, and recommendations within this article are solely those of the authors. The article authors have declared no conflicts of interest.